OpenStudy (anonymous):

Help please solve the equation 3(c-2)=2(c-6) 2 -6 No Solution Identity

OpenStudy (anonymous):

Distribute (meaning multiply) the numbers on outside of paranthesis to each number/letter on inside.. 3(c-2)=2(c-6) 3c-6=2c-12 Now get letters on one side and numbers on another subtract by 2c to both sides c-6=-12 (now add 6 to both sides) c=-6 I hope that helps

OpenStudy (anonymous):

yep, I agree with the answer too that is what I got.